OpenStudy (gabbyalicorn):

Uri asked six friends how many different foreign countries they’d been to. He recorded 0, 1, 1, 2, 5, and 9 countries visited. What is the mean absolute deviation of the number of countries visited? Round to the nearest tenth, if needed. A. 0 countries B. 2.7 countries C. 2.9 countries D. 3 countries same thing? @black_ninja123 ?

OpenStudy (black_ninja123):

Yes

OpenStudy (gabbyalicorn):

i got 3 for step 1

OpenStudy (gabbyalicorn):

then 0-3= 3 1-3 = 2 1-3 = 2 2-3 = 1 5-3 = 2 9-3 = 6

OpenStudy (black_ninja123):

Yes

OpenStudy (gabbyalicorn):

i got 2.6 which would be B asuming that it got rounded?

OpenStudy (black_ninja123):

Correct

OpenStudy (gabbyalicorn):

:D